# Fast data inversion for high-dimensional dynamical systems from noisy measurements

### v2

@@ 2.2 A fast EM algorithm with closed-form expressions
In this section, we derive a fast EM algorithm for a general scenario where the parameters are $\mathbf{\Theta}=(\mathbf{U}_{0},\sigma_{0}^{2},\bm{\rho},\bm{\sigma}^{2})$ with $\bm{\rho}=(\rho_{1},\rho_{2},\dots,\rho_{d})$ and $\bm{\sigma}^{2}=(\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{d}^{2})$. The algorithm leverages KF and RTS smoother for computing required quantities with orthogonal projection, to bypass costly matrix inversion, conventionally required in each step of the KF. A short review of the KF and RTS smoother is provided in Section S2 and the proofs in this subsection are given in Section S3 in the supplementary materials. We call this approach a Fast algorithm of Multivariate Ornstein-Uhlenbeck processes (FMOU).
@@
Direct optimization of the marginal likelihood of the model in Equations (3) and (4) can be unstable due to optimizing the latent factor matrix in the Stiefel manifold in each step of the optimization (Wen and Yin,, 2013) and numerical optimization in high-dimensional parameter space. We develop an EM algorithm which has closed-form expressions in each iteration.
@@
$\ell(\bm{\Theta})=C-\frac{nk}{2}\log(\sigma^{2}_{0})-\frac{\operatorname{tr}(\mathbf{Y}^{T}\mathbf{Y}-2\mathbf{Y}^{T}\mathbf{U}_{0}\mathbf{Z})}{2\sigma^{2}_{0}}-\sum^{d}_{l=1}\Bigl{(}\frac{\mathbf{z}_{l}^{T}\mathbf{z}_{l}}{2\sigma^{2}_{0}}+\frac{\log|\bm{\Sigma}_{l}|+\mathbf{z}_{l}^{T}\bm{\Sigma}^{-1}_{l}\mathbf{z}_{l}}{2}\Bigr{)},$ (9)
@@
In the E step, we calculate the expectation of the joint log-likelihood function with respect to the distribution of $\mathbf{Z}$ conditional on observations $\mathbf{Y}$, and the current estimate of parameters, denoted as $\bm{\hat{\Theta}}=(\mathbf{\hat{U}}_{0},\hat{\sigma}_{0}^{2},\bm{\hat{\rho}},\bm{\hat{\sigma}}^{2})$. The posterior distribution of factors given $\bm{\hat{\Theta}}$ follows a multivariate Gaussian distribution (Gu and Shen,, 2020):
$(\mathbf{z}_{l}\mid\mathbf{Y},\bm{\hat{\Theta}})\sim\mathcal{MN}\left(\mathbf{\hat{z}}_{l},\hat{\sigma}^{2}_{0}\bm{\hat{\Sigma}}_{l}({\bm{\hat{\Sigma}}}_{l}+\hat{\sigma}^{2}_{0}\mathbf{I}_{n})^{-1}\right),$ (10)
@@
$\quad-\sum^{d}_{l=1}\frac{\mathbf{\hat{z}}_{l}^{T}\bm{\Sigma}^{-1}_{l}\mathbf{\hat{z}}_{l}+\mbox{tr}[{\hat{\sigma}}_{0}^{2}\bm{\Sigma}^{-1}_{l}\bm{\hat{\Sigma}}_{l}(\bm{\hat{\Sigma}}_{l}+{\hat{\sigma}}_{0}^{2}\mathbf{I}_{n})^{-1}]}{2}-\frac{nk}{2}\log(\sigma^{2}_{0})-\frac{\operatorname{tr}(\mathbf{Y}^{T}\mathbf{Y})}{2\sigma^{2}_{0}},$ (11)
where $\bm{\Sigma}_{l}$ depends on unknown $(\bm{\rho},\bm{\sigma}^{2})$ to be optimized, whereas $\bm{\hat{\Sigma}}_{l}$ is obtained based on the current estimate of the parameters $(\bm{\hat{\rho}},\hat{\bm{\sigma}}^{2})$.
@@
Consider a dynamic linear model with output $\tilde{y}_{l}(t)=z_{l}(t)+\epsilon$ and latent factor process defined in (4), with $\epsilon$ being an independent noise with variance $\sigma^{2}_{0}$ for $l=1,...,d$. Denote $s_{l}(t)=\operatorname{\mathbb{E}}[z_{l}(t)\mid\mathbf{\tilde{y}}_{l},\hat{\bm{\Theta}}]$, $S_{l}(t)=\operatorname{\mathbb{V}}[z_{l}(t)\mid\mathbf{\tilde{y}}_{l},\hat{\bm{\Theta}}]$ and $\tilde{S}_{l}(t)=\operatorname{Cov}[z_{l}(t),z_{l}(t+1)\mid\mathbf{\tilde{y}}_{l},\hat{\bm{\Theta}}]$, which are computed by the KF and RTS smoother. By leveraging the tridiagonal structure of $\mathbf{\Sigma}_{l}^{-1}$, we have
$\operatorname{tr}[{\hat{\sigma}}_{0}^{2}\bm{\hat{\Sigma}}_{l}(\bm{\hat{\Sigma}}_{l}+{\hat{\sigma}}_{0}^{2}\mathbf{I}_{n})^{-1}]=\sum_{t=1}^{n}\operatorname{\mathbb{V}}[\mathbf{z}_{l}(t)\mid\mathbf{Y},\hat{\bm{\Theta}}]=\sum_{t=1}^{n}S_{l}(t),$ (12)
$\mathbf{\hat{z}}_{l}^{T}\mathbf{\Sigma}_{l}^{-1}\mathbf{\hat{z}}_{l}=\frac{1}{\sigma_{l}^{2}}\left((1-\rho_{l}^{2})s_{l}(1)+\sum_{t=2}^{n}(s_{l}(t)-\rho_{l}s_{l}(t-1))^{2}\right),$ (13)
$\operatorname{tr}[{\hat{\sigma}}_{0}^{2}\bm{\Sigma}^{-1}_{l}\bm{\hat{\Sigma}}_{l}(\bm{\hat{\Sigma}}_{l}+{\hat{\sigma}}_{0}^{2}\mathbf{I}_{n})^{-1}]=\frac{1}{\sigma_{l}^{2}}\left(\sum_{t=1}^{n}S_{l}(t)+\rho_{l}^{2}\sum_{t=2}^{n-1}S_{l}(t)-2\rho_{l}\sum_{t=1}^{n-1}\tilde{S}_{l}(t)\right).$ (14)
@@
$\bm{\hat{\Theta}}^{\text{new}}=\mbox{argmax}_{\bm{\Theta}}\bar{\ell}(\bm{\Theta}).$ (15)
We emphasize that all parameters have closed-form expressions in the M-step in (15), thus avoiding numerically optimizing the high-dimensional parameter space in the algorithm. An efficient update in M-step is given in Theorem 2.5.
@@
    $\hat{\mathbf{U}}_{0}^{\text{new}}=\tilde{\mathbf{V}}\tilde{\mathbf{U}}^{T},$ (16)
@@
    $({\hat{\sigma}}^{2}_{0})^{new}=\frac{\operatorname{tr}(\mathbf{Y}^{T}\mathbf{Y})-2\operatorname{tr}(\mathbf{Y}^{T}\mathbf{\hat{U}}_{0}^{\text{new}}\hat{\mathbf{Z}})+\sum_{l=1}^{d}\sum_{t=1}^{n}(s^{2}_{l}(t)+S_{l}(t))}{nk}.$ (17)
@@
    $\beta_{0}+\beta_{1}\rho_{l}+\beta_{2}\rho_{l}^{2}+\beta_{3}\rho_{l}^{3}=0,$ (18)
@@
    $\qquad\frac{1}{n}\left\{\sum_{t=1}^{n}S_{l}(t)+(\hat{\rho}_{l}^{\text{new}})^{2}\sum_{t=2}^{n-1}S_{l}(t)-2\hat{\rho}_{l}^{\text{new}}\sum_{t=1}^{n-1}\tilde{S}_{l}(t)\right\}.$ (19)
@@
For our application of slip estimation from GPS data, the latent loading matrix is computed by Green’s function, and the variance of the noise is usually available from GPS measurement in practice. The FMOU is applicable to this scenario as well, discussed in Section S3.5 in the supplementary materials. Algorithm 1 summarizes our FMOU approach.
The output of Algorithm 1 is the MMLE of the parameters $\bm{\Theta}^{\text{MMLE}}$ in (8) and posterior mean of the latent factors given $\bm{\Theta}^{\text{MMLE}}$, denoted as $\mathbf{Z}^{\text{post}}=[\mathbf{z}^{\text{post}}(1),...,\mathbf{z}^{\text{post}}(n)]=\mathbb{E}[\mathbf{Z}\mid\mathbf{Y},\bm{\Theta}^{\text{MMLE}}]$. In our application of slip estimation, other quantities, such as the posterior distribution of the mean of the observations and slips at time $t$ follows
$\left(\mathbf{m}(t)\mid\mathbf{Y},{\bm{\Theta}}^{\text{MMLE}}\right)\sim\mathcal{MN}\left({\mathbf{U}}_{0}{\mathbf{z}}^{\text{post}}(t),{\mathbf{\Sigma}}_{m}^{\text{MMLE}}(t)\right),$ (20)
$\left(\mathbf{z}_{s}(t)\mid\mathbf{Y},{\bm{\Theta}}^{\text{MMLE}}\right)\sim\mathcal{MN}\left(\mathbf{G}^{T}{\mathbf{U}}_{0}\mathbf{D}_{0}^{-2}{\mathbf{z}}^{\text{post}}(t),{\mathbf{\Sigma}}^{\text{MMLE}}_{s}(t)\right),$ (21)
where ${\mathbf{\Sigma}}^{\text{MMLE}}_{m}(t)={\mathbf{U}}_{0}\mathbf{D}_{S}(t){\mathbf{U}}_{0}^{T}$ and $\mathbf{\Sigma}^{\text{MMLE}}_{s}(t)=\mathbf{G}^{T}{\mathbf{U}}_{0}\mathbf{D}_{0}^{-2}\mathbf{D}_{S}(t)\mathbf{D}_{0}^{-2}{\mathbf{U}}_{0}^{T}\mathbf{G}$, with $\mathbf{D}_{S}(t)$ being a $d\times d$ diagonal matrix where the $l$th element is $S_{l}(t)$ computed by plugging in the MMLE of the parameters for $l=1,2,\dots,d$. Here $\mathbf{U}_{0}$ can be either fixed or estimated. In the application of slip estimation, $\mathbf{U}_{0}$ and $\mathbf{D}_{0}$ are the first $d$ left singular vectors and the largest $d$ singular values of the matrix of Green’s function $\mathbf{G}$, respectively.
@@
Several advantages of this EM algorithm are worth mentioning. First, the kernel parameters in Gaussian processes are notoriously hard to estimate numerically, and we have $d$ sets of kernel and scale parameters, where $d$ can be on the order of $10^{3}$ or even $10^{6}$ in practice. The estimates of these parameters have closed-form expressions in our EM algorithm, which had not been derived before. Our algorithm substantially improves estimation stability compared to optimization algorithms that numerically optimize in a high-dimensional parameter space. Second, when $\mathbf{U}_{0}$ is unknown and the parameters are distinct in each latent process, originally one needs to solve a constrained optimization problem in the Stiefel manifold (Wen and Yin,, 2013) to optimize $\mathbf{U}_{0}$ in each of the numerical iteration for estimating the parameters of the model in (3) and (4) (Gu and Shen,, 2020). In contrast, we derived closed-form expressions for estimating $\mathbf{U}_{0}$ in each iteration of the EM algorithm, and hence the algorithm is much faster. Third, we integrate orthogonal projection in KF and RTS smoother to bypass the matrix inversion operation at each time step. For $M$ iterations of the EM algorithms, the overall complexity of the parameter estimation process is $\mathcal{O}(Mknd)+\mathcal{O}(Mkd^{2})$ when the coefficient $\mathbf{U}_{0}$ is unknown. For the scenario where $\mathbf{U}_{0}$ is derived from the Green’s function, we need to perform SVD to obtain $\mathbf{U}_{0}$, which has the order $\mathcal{O}(min({k}^{2}k^{\prime},{k}(k^{\prime})^{2}))$.

### v3

@@ 2.3 A fast EM algorithm with closed-form expressions
In this section, we derive a fast EM algorithm, named a Fast algorithm of Multivariate Ornstein-Uhlenbeck processes (FMOU), for a general scenario where the parameters are $\mathbf{\Theta}=(\mathbf{U}_{0},\sigma_{0}^{2},\bm{\rho},\bm{\sigma}^{2})$ with $\bm{\rho}=(\rho_{1},\rho_{2},\dots,\rho_{d})$ and $\bm{\sigma}^{2}=(\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{d}^{2})$. The algorithm leverages KF [21] and RTS smoother [30] for computing required quantities with orthogonal projection, to bypass costly matrix inversion, conventionally required in each step of the KF. A short review of the KF and RTS smoother [43, 28] is provided in Section SM2. The proofs in this subsection are given in Section SM3 in the supplementary materials.
@@
Direct optimization of the marginal likelihood of the model in Equations (3) and (4) can be unstable due to optimizing the latent factor matrix in the Stiefel manifold in each step of the optimization [42] and numerical optimization in high-dimensional parameter space. We develop an EM algorithm which has closed-form expressions in each iteration.
@@
(9) $\ell(\bm{\Theta})=C-\frac{nk}{2}\log(\sigma^{2}_{0})-\frac{\operatorname{tr}(\mathbf{Y}^{T}\mathbf{Y}-2\mathbf{Y}^{T}\mathbf{U}_{0}\mathbf{Z})}{2\sigma^{2}_{0}}-\sum^{d}_{l=1}\Bigl{(}\frac{\mathbf{z}_{l}^{T}\mathbf{z}_{l}}{2\sigma^{2}_{0}}+\frac{\log|\bm{\Sigma}_{l}|+\mathbf{z}_{l}^{T}\bm{\Sigma}^{-1}_{l}\mathbf{z}_{l}}{2}\Bigr{)},$
@@
In the E step, we calculate the expectation of the joint log-likelihood function with respect to the distribution of $\mathbf{Z}$ conditional on observations $\mathbf{Y}$, and the current estimate of parameters, denoted as $\bm{\hat{\Theta}}=(\mathbf{\hat{U}}_{0},\hat{\sigma}_{0}^{2},\bm{\hat{\rho}},\bm{\hat{\sigma}}^{2})$. The posterior distribution of factors given $\bm{\hat{\Theta}}$ follows a multivariate Gaussian distribution [18]:
(10) $(\mathbf{z}_{l}\mid\mathbf{Y},\bm{\hat{\Theta}})\sim\mathcal{MN}\left(\mathbf{\hat{z}}_{l},\hat{\sigma}^{2}_{0}\bm{\hat{\Sigma}}_{l}({\bm{\hat{\Sigma}}}_{l}+\hat{\sigma}^{2}_{0}\mathbf{I}_{n})^{-1}\right),$
@@
(11) $\quad-\sum^{d}_{l=1}\frac{\mathbf{\hat{z}}_{l}^{T}\bm{\Sigma}^{-1}_{l}\mathbf{\hat{z}}_{l}+\mbox{tr}[{\hat{\sigma}}_{0}^{2}\bm{\Sigma}^{-1}_{l}\bm{\hat{\Sigma}}_{l}(\bm{\hat{\Sigma}}_{l}+{\hat{\sigma}}_{0}^{2}\mathbf{I}_{n})^{-1}]}{2}-\frac{nk}{2}\log(\sigma^{2}_{0})-\frac{\operatorname{tr}(\mathbf{Y}^{T}\mathbf{Y})}{2\sigma^{2}_{0}},$
where $\bm{\Sigma}_{l}$ depends on unknown $(\bm{\rho},\bm{\sigma}^{2})$ to be optimized, and $\bm{\hat{\Sigma}}_{l}$ is obtained based on the current estimate of the parameters $(\bm{\hat{\rho}},\hat{\bm{\sigma}}^{2})$.
@@
Consider a dynamic linear model $\tilde{y}_{l}(t)=z_{l}(t)+\epsilon$ with latent factor process defined in (4) and $\epsilon$ being an independent noise with variance $\sigma^{2}_{0}$, for $l=1,...,d$. Denote $s_{l}(t)=\operatorname{\mathbb{E}}[z_{l}(t)\mid\mathbf{\tilde{y}}_{l},\hat{\bm{\Theta}}]$, $S_{l}(t)=\operatorname{\mathbb{V}}[z_{l}(t)\mid\mathbf{\tilde{y}}_{l},\hat{\bm{\Theta}}]$ and $\tilde{S}_{l}(t)=\operatorname{Cov}[z_{l}(t),z_{l}(t+1)\mid\mathbf{\tilde{y}}_{l},\hat{\bm{\Theta}}]$, which are computed by the KF and RTS smoother. We have
(12) $\operatorname{tr}[{\hat{\sigma}}_{0}^{2}\bm{\hat{\Sigma}}_{l}(\bm{\hat{\Sigma}}_{l}+{\hat{\sigma}}_{0}^{2}\mathbf{I}_{n})^{-1}]=\sum_{t=1}^{n}\operatorname{\mathbb{V}}[\mathbf{z}_{l}(t)\mid\mathbf{Y},\hat{\bm{\Theta}}]=\sum_{t=1}^{n}S_{l}(t),$
(13) $\mathbf{\hat{z}}_{l}^{T}\mathbf{\Sigma}_{l}^{-1}\mathbf{\hat{z}}_{l}=\frac{1}{\sigma_{l}^{2}}\left((1-\rho_{l}^{2})s_{l}(1)+\sum_{t=2}^{n}(s_{l}(t)-\rho_{l}s_{l}(t-1))^{2}\right),$
(14) $\operatorname{tr}[{\hat{\sigma}}_{0}^{2}\bm{\Sigma}^{-1}_{l}\bm{\hat{\Sigma}}_{l}(\bm{\hat{\Sigma}}_{l}+{\hat{\sigma}}_{0}^{2}\mathbf{I}_{n})^{-1}]=\frac{1}{\sigma_{l}^{2}}\left(\sum_{t=1}^{n}S_{l}(t)+\rho_{l}^{2}\sum_{t=2}^{n-1}S_{l}(t)-2\rho_{l}\sum_{t=1}^{n-1}\tilde{S}_{l}(t)\right).$
@@
(15) $\bm{\hat{\Theta}}^{\text{new}}=\mbox{argmax}_{\bm{\Theta}}\bar{\ell}(\bm{\Theta}).$
In Theorem 2.5, we show that all parameters have closed-form expressions in the M-step in (15), thus avoiding numerically optimizing high-dimensional parameters in the algorithm.
@@
    (16) $\hat{\mathbf{U}}_{0}^{\text{new}}=\tilde{\mathbf{V}}\tilde{\mathbf{U}}^{T},$
@@
    (17) $({\hat{\sigma}}^{2}_{0})^{new}=\frac{\operatorname{tr}(\mathbf{Y}^{T}\mathbf{Y})-2\operatorname{tr}(\mathbf{Y}^{T}\mathbf{\hat{U}}_{0}^{\text{new}}\hat{\mathbf{Z}})+\sum_{l=1}^{d}\sum_{t=1}^{n}(s^{2}_{l}(t)+S_{l}(t))}{nk}.$
@@
    (18) $\beta_{0}+\beta_{1}\rho_{l}+\beta_{2}\rho_{l}^{2}+\beta_{3}\rho_{l}^{3}=0,$
@@
    (19) $\qquad\frac{1}{n}\left\{\sum_{t=1}^{n}S_{l}(t)+(\hat{\rho}_{l}^{\text{new}})^{2}\sum_{t=2}^{n-1}S_{l}(t)-2\hat{\rho}_{l}^{\text{new}}\sum_{t=1}^{n-1}\tilde{S}_{l}(t)\right\}.$
@@
The FMOU in Theorem 2.5 is nontrivial to derive, and it differs from the conventional EM algorithm for vector autoregressive (VAR) models [34]. First, FMOU avoids the costly matrix inversion required in each step in the KF due to the use of an orthogonal latent factor loading matrix that can be estimated from data. Second, it enables dimension reduction. Third, the FMOU algorithm can be set to be stationary and nonstationary, while the conventional EM for VAR models cannot guarantee stationarity in estimation.
Algorithm 1 summarizes our FMOU approach. The output of Algorithm 1 is the MMLE of the parameters $\bm{\Theta}^{\text{MMLE}}$ in (8) and posterior mean of the latent factors given $\bm{\Theta}^{\text{MMLE}}$, denoted as $\mathbf{Z}^{\text{post}}=[\mathbf{z}^{\text{post}}(1),...,\mathbf{z}^{\text{post}}(n)]=\mathbb{E}[\mathbf{Z}\mid\mathbf{Y},\bm{\Theta}^{\text{MMLE}}]$. In our application of slip estimation, other quantities, such as the posterior distribution of the mean of the observations and slips at time $t$ follows
(20) $\left(\mathbf{m}(t)\mid\mathbf{Y},{\bm{\Theta}}^{\text{MMLE}}\right)\sim\mathcal{MN}\left({\mathbf{U}}_{0}{\mathbf{z}}^{\text{post}}(t),{\mathbf{\Sigma}}_{m}^{\text{MMLE}}(t)\right),$
(21) $\left(\mathbf{z}_{s}(t)\mid\mathbf{Y},{\bm{\Theta}}^{\text{MMLE}}\right)\sim\mathcal{MN}\left(\mathbf{G}^{T}{\mathbf{U}}_{0}\mathbf{D}_{0}^{-2}{\mathbf{z}}^{\text{post}}(t),{\mathbf{\Sigma}}^{\text{MMLE}}_{s}(t)\right),$
where ${\mathbf{\Sigma}}^{\text{MMLE}}_{m}(t)={\mathbf{U}}_{0}\mathbf{D}_{S}(t){\mathbf{U}}_{0}^{T}$ and $\mathbf{\Sigma}^{\text{MMLE}}_{s}(t)=\mathbf{G}^{T}{\mathbf{U}}_{0}\mathbf{D}_{0}^{-2}\mathbf{D}_{S}(t)\mathbf{D}_{0}^{-2}{\mathbf{U}}_{0}^{T}\mathbf{G}$, with $\mathbf{D}_{S}(t)$ being a $d\times d$ diagonal matrix where the $l$th element is $S_{l}(t)$ computed by plugging in the MMLE of the parameters for $l=1,2,\dots,d$. Here $\mathbf{U}_{0}$ can be either fixed or estimated. In the application of slip estimation, $\mathbf{U}_{0}$ and $\mathbf{D}_{0}$ are the first $d$ left singular vectors and the largest $d$ singular values of the matrix of Green’s function $\mathbf{G}$, respectively and variance of the noise is usually available from GPS measurements, with details discussed in Section SM3.5. Furthermore, the initialization of $\mathbf{U}_{0}$ and $\sigma^{2}_{0}$ can be initialized as an identity matrix and uniform distribution, respectively, as they do not have a large impact on estimating the mean due to the flexibility of the models from a large number of parameters.
@@
Several advantages of this EM algorithm are worth mentioning. First, the kernel parameters in Gaussian processes are notoriously hard to estimate numerically, and we have $d$ sets of kernel and scale parameters, where $d$ can be on the order of $10^{3}$ or even $10^{6}$ in some applications. The estimates of these parameters have closed-form expressions in our EM algorithm, which had not been derived before. Our algorithm substantially improves estimation stability compared to optimization algorithms that numerically optimize in a high-dimensional parameter space. Second, when $\mathbf{U}_{0}$ is unknown and the parameters are distinct in each latent process, originally one needs to solve a constrained optimization problem in the Stiefel manifold [42] to optimize $\mathbf{U}_{0}$ in each of the numerical iteration for estimating the parameters of the model in (3) and (4) [18]. In contrast, we derived closed-form expressions for estimating $\mathbf{U}_{0}$ in each iteration of the EM algorithm, and hence the algorithm is much faster. Third, we integrate orthogonal projection in the KF and RTS smoother to bypass the matrix inversion operation at each time step. For $M$ iterations of the EM algorithms, the overall complexity of the parameter estimation process is $\mathcal{O}(Mknd)+\mathcal{O}(Mkd^{2})$ when the coefficient $\mathbf{U}_{0}$ is unknown, as shown in Table 1. For the scenario where $\mathbf{U}_{0}$ is derived from the Green’s function, we need to perform SVD to obtain $\mathbf{U}_{0}$ only once, which has the order $\mathcal{O}(min({k}^{2}k^{\prime},{k}(k^{\prime})^{2}))$.